# Tunable non-Markovian dynamics in a collision model: an application to coherent transport

### v1

@@ Abstract
**footnotetext: These authors contributed equally to this work
@@ I Introduction
The study of coherent energy transport in quantum systems is of great theoretical and technological importance in physics, chemistry, biology and engineering [1, 2, 3, 4, 5, 6]. In photosynthetic systems, long-lived quantum coherence was observed in excitonic energy transfer [7, 8], but its role in the remarkable efficiency of the process has been debated [9, 10, 11]. Moreover, the role of the environment in this process has been shown to be complex [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
@@
Nevertheless, photosynthetic systems provide an ideal scenario where recently proposed witnesses of non-classicality could be tested. These witnesses suggest the possibility of indirectly inferring the non-classicality of an unknown system by looking at its capability of creating quantum coherence in a quantum probe [24]. For example, the non-classicality of a generic biopolymer could be inferred by observing the coherence of a photon emitted through the recombination of a delocalised exciton on the polymer [25]. To investigate the possibility of this task, it is essential to understand the mechanism for coherent transport in these systems and clarify the role of the environment in the process.
@@
In this paper, we take an information-theoretic approach to the problem. We propose a collision model [26, 27] to study the non-Markovian information dynamics of a chain of qubits interacting with the environment. Differently from other models in the literature [28, 29, 30, 31, 32, 33, 34, 35], the system interacts locally with the same reservoir qubits and we control the degree of non-Markovianity of the environment by applying a depolarising channel to the reservoir. By changing the intensity of the depolarising effect, we can control the amount of information irretrievably lost by the environment.
@@
We first quantify the degree of non-Markovianity that the depolarising channel induces in the system. Then, we use the collision model to study the coherent transport of an excitation along a chain of qubits coupled to a fixed and rigid reservoir of interacting qubits. By changing the degree of non-Markovianity of the reservoir, we can model different types of environments. For example, in an analogy to the Fenna-Matthews-Olson (FMO) photosynthetic complex [36], a Markovian environment could represent the solvent surrounding the complex, while a strongly non-Markovian one could model the phonons of the bacteriochlorophyll molecules that couple to the electronic degrees of freedom responsible for the excitonic transfer.
@@
We show that the transfer of information to the environment plays a crucial role in the coherent transfer of the excitation. The process displays different behaviours depending on the strength of the system-environment coupling. When the system-environment coupling is strong, non-Markovianity dramatically improves the efficiency of coherent transport. Interestingly, when the coupling is weak, a Markovian environment is preferable to enhance the amount of coherence and the coherence time of the excitation.
@@
Our results provide some general insights into how non-Markovian environments affect coherent excitonic transport on a chain. The generality of this model makes it valuable in several fields where energy transport phenomena are relevant. The model also demonstrates how to introduce a control on the degree of non-Markovianity in collision models with a fixed and rigid environment.
@@ III.2 Application to coherent transport
We apply the model described above to study the coherent transport of an excitation along the chain. We represent an excitation localised on the first qubit as
@@
in the Hilbert space $\mathscr{H}_{S_{1}}\otimes\dots\otimes\mathscr{H}_{S_{N}}$. Applying the protocol described above, the excitation spreads through the chain and the environment. After repeating the protocol $k$ times we trace out the environment and extract the coherence between the first and last qubit of the chain $\rho_{C}^{1,N}$. We take $|\rho_{C}^{1,N}|$ as an indicator of the coherent transport efficiency across the chain.
Figure 2: Quantum circuit implementing an iteration of the protocol in our model. The three pswap gates implement the exchange phase. A depolarising error is associated with each identity gate to implement the depolarisation phase. The two final gates implement the unitary evolutions with the Hamiltonians $\hat{H}_{res}$ and $\hat{H}_{chain}$ of the transfer phases.
For simplicity, we consider a model with three chain qubits. Fig. 2 shows how an iteration of the protocol can be implemented as a quantum circuit in Qiskit [39]. The exchange phase consists of a sequence of local unitaries $\hat{P}(\eta)$ as in Eq. (4) for each of the system qubits $S_{i}$; the depolarisation phase is implemented as a depolarising noise model on the identity gate, which applies the quantum channel of Eq. (6); the transfer phases are implemented as unitaries on system and reservoir with a time step $\Delta t=0.01$. We choose $J^{n}_{chain}=J_{chain}=10,\,\forall n$ and $J^{m}_{res}=J_{res}=1,\,\forall m$. This choice represents a scenario where the chain is the most efficient medium for the propagation of the excitation.
@@ IV.2 Coherent transport
Overall, we can conclude that non-Markovianity improves the coherent transport for strong system-environment couplings but hinders it if the coupling is weak. In the weak coupling regime, a Markovian environment enhances both the amount of coherence and the coherence time of the excitation. Note that these two mechanisms may not be mutually exclusive. For example, in the FMO photosynthetic complex, the electronic degrees of freedom of the bacteriochlorophyll molecules (the chain in our model) can couple both to the phonons of the molecular environment (non-Markovian) and to an external thermal bath (Markovian environment). Our results capture both of these situations, suggesting that the coexistence of a weakly coupled thermal bath and a strongly coupled non-Markovian environment may determine an optimal condition for coherent excitonic transport.
@@ V Conclusions
We have applied our model to study the quantum coherent excitonic transport on a chain of interacting qubits coupled to an environment. Our model permits the exploration of how different regimes of non-Markovianity and system-environment coupling affect coherent transport. Studying this aspect is important in photosynthetic complexes, where the electronic degrees of freedom responsible for the coherent transport of an exciton can couple to both Markovian (e.g. solvent) and non-Markovian (phonons) environments. Our results show that a non-Markovian environment aids coherent transport only in the strong system-environment coupling regime, while a Markovian environment is surprisingly preferable in the weak coupling regime.
This work provides some general insights into how non-Markovianity affects coherent excitonic transport, avoiding the complications that arise from accurately modelling a real complex physical system. Its generality makes it suitable for applications to any quantum processes whose understanding involves the exchange of information between the system and the environment. The model could also be easily adapted to study more specific systems. For example, one could expand the dimensions of the chain and the environment, use more realistic Hamiltonians, or use two different environments, one Markovian and one non-Markovian. These generalisations could be relevant to inform the development of novel technological applications. The model could also be implemented on current quantum computer [40]. We leave these investigations for future work.

### v2

@@ Abstract
**footnotetext: These authors contributed equally to this work.
@@ I Introduction
The study of coherent energy transport in quantum systems is of great theoretical and technological importance in physics, chemistry, biology and engineering [1, 2, 3, 4, 5, 6]. In photosynthetic systems, long-lived quantum coherence was observed in excitonic energy transfer [7, 8], although its role in the remarkable efficiency of the process has been debated [9, 10, 11]. Moreover, the role of the environment in this process has been shown to be complex [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Nevertheless, photosynthetic systems provide a promising scenario in which recently proposed witnesses of non-classicality could be tested. These witnesses suggest the possibility of indirectly inferring the non-classicality of an unknown system by looking at its capability of creating quantum coherence in a quantum probe [25]. For example, the non-classicality of a generic biopolymer could be inferred by observing the coherence of a photon emitted through the recombination of a delocalised exciton on the polymer [26]. To investigate the possibility of this task, it is essential to understand the mechanism for coherent transport in these systems and clarify the role of the environment in the process.
@@
In this paper, we take an information-theoretic approach to investigate coherent transport in a simplified scenario. We propose a collision model [27, 28] to study the non-Markovian information dynamics of a chain of qubits interacting with the environment. Differently from other models in the literature [29, 30, 31, 32, 33, 34, 35, 36], the system interacts locally with the same reservoir qubits and we control the degree of non-Markovianity of the environment by applying a depolarising channel to the reservoir. By changing the intensity of the depolarising effect, we can control the amount of information irretrievably lost by the environment.
@@
We first quantify the degree of non-Markovianity that the depolarising channel induces in the system. Then, we use the collision model to study the coherent transport of an excitation along a chain of qubits coupled to a fixed and rigid reservoir of interacting qubits. By changing the degree of non-Markovianity of the reservoir, we can model different types of environments. For example, a Markovian environment could represent a thermal bath surrounding the chain, while a non-Markovian one could model other degrees of freedom of the chain that are coupled to the ones responsible for the coherent transport.
@@
We show that the transfer of information to the environment plays a crucial role in the coherence of the excitonic transport. The process displays different behaviours depending on the strength of the system-environment coupling. When the system-environment coupling is strong, non-Markovianity dramatically improves the coherence of the transport. Interestingly, when the coupling is weak, a Markovian environment is preferable to enhance the amount of coherence and the coherence time of the excitation.
@@
Our results provide some information-theoretic insights into how non-Markovian environments affect the coherence of the excitonic transport on a chain. Moreover, our model demonstrates how to introduce a control on the degree of non-Markovianity in collision models with a fixed environment and might be applied in other fields where energy transport phenomena are relevant.
@@ III.2 Application to coherent transport
We apply the model described above to study the coherence of the transport of an excitation along the chain. We represent an excitation localised on the first qubit as
@@
in the Hilbert space $\mathscr{H}_{S_{1}}\otimes\dots\otimes\mathscr{H}_{S_{N}}$. Applying the protocol described above, the excitation spreads through the chain and the environment. After repeating the protocol $k$ times, we trace out the environment and extract the coherence between the first and last qubit of the chain $\rho_{C}^{1,N}$, that is, the coefficient of the term |$\ket{100\dots 0}\bra{00\dots 01}$ | of the density matrix. We take $|\rho_{C}^{1,N}|$ as an indicator of the coherent transport efficiency across the chain.
For simplicity, we consider a model with three chain qubits. We implemented the protocol as a quantum circuit in Qiskit [39] as follows. The exchange phase consists of a sequence of local unitaries $\hat{P}(\eta)$ as in Eq. (4) for each of the system qubits $S_{i}$; the depolarisation phase is implemented as a depolarising noise model on the identity gate, which applies the quantum channel of Eq. (6); the transfer phases are implemented as unitaries on system and reservoir with a time step $\Delta t=0.01$. We choose $J^{n}_{chain}=J_{chain}=10,\,\forall n$ and $J^{m}_{res}=J_{res}=1,\,\forall m$. This choice represents a scenario where the chain is the most efficient medium for the propagation of the excitation.
@@ IV.2 Coherent transport
Overall, we can conclude that non-Markovianity improves the coherent transport for strong system-environment couplings but hinders it if the coupling is weak. In the weak coupling regime, a Markovian environment enhances both the amount of coherence and the coherence time of the excitation. Note that these two mechanisms may not be mutually exclusive. For example, in the Fenna-Matthews-Olson photosynthetic complex [40], the electronic degrees of freedom of the bacteriochlorophyll molecules (the chain in our model) can couple both to the phonons of the molecular environment (non-Markovian) and to an external thermal bath (Markovian environment). Our results might capture both of these situations, suggesting that the coexistence of a weakly coupled thermal bath and a strongly coupled non-Markovian environment might determine an optimal condition for coherent excitonic transport.
@@ V Conclusions
We have applied our model to study the quantum coherent excitonic transport on a chain of interacting qubits coupled to an environment. Our model permits the exploration of how different regimes of non-Markovianity and system-environment coupling affect coherent transport. Studying this aspect is important, for instance, in photosynthetic complexes, where the electronic degrees of freedom responsible for the coherent transport of an exciton can couple to both Markovian (e.g. solvent) and non-Markovian (phonons) environments. Our results show that a non-Markovian environment aids coherent transport only in the strong system-environment coupling regime, while a Markovian environment is surprisingly preferable in the weak coupling regime.
This work provides some information-theoretic insights into how non-Markovianity affects coherent excitonic transport, avoiding the complications that arise from accurately modelling a real complex physical system. The system-independent characteristic of the model might make it suitable for applications to any quantum processes whose understanding involves the exchange of information between the system and the environment. The model could also be easily adapted to study more specific systems. For example, one could expand the dimensions of the chain and the environment, use more realistic Hamiltonians, or use two different environments, one Markovian and one non-Markovian. These generalisations might be relevant to inform the development of novel technological applications. The model could also be implemented on current quantum computers [41]. We leave these investigations for future work.